OpenStudy (anonymous):

1. Find the first six terms of the sequence. a1 = -8, an = 5 • an-1 -8, -40, -200, -1000, -5000, -25,000 -8, -40, -35, -30, -25, -20 0, 5, -40, -35, -30, -25 -40, -200, -1000, -5000, -25,000, -125,000 2.) Find an equation for the nth term of the arithmetic sequence. a(subscript)19 = -58, a(subscript)21 = -164 an = 896 - 53(n - 2) an = 896 - 53(n - 1) an = 896 + 53(n + 1) an = 896 - 53(n + 1) 3. Find an equation for the nth term of the arithmetic sequence. -1, 2, 5, 8, ... an = -1 + 3(n) an = -1 + 3(n - 1) an = -1 + 3 an = -1 + 3(n + 1)

OpenStudy (princeharryyy):

q1.

OpenStudy (princeharryyy):

a2 = 5.a1 = 5*-8 = -40 a3 = 5*a2 = 5*-40 = -200 a4 = 5*-200 = -1000 and so on. So, the answer is a @em2000

OpenStudy (anonymous):

got it @princeharryyy

OpenStudy (princeharryyy):

q2.

OpenStudy (princeharryyy):

a19 = a + 18d = -58. let it be 1 a21 = a + 20d = -168 let it be 2 equation(2-1) a+20d - a-18d =-164 +58 2d = -106 d = -53 putting in 1 a=-58+18*53 = 954-58 = 896

OpenStudy (princeharryyy):

so answer => 896 -53(n-1) @em2000

OpenStudy (princeharryyy):

that's is b. r u getting it. @em2000

OpenStudy (princeharryyy):

an for any ap is given as an = a + (n-1)*d is known as general term @em2000 u getting.

OpenStudy (princeharryyy):

q.3

OpenStudy (princeharryyy):

-1, 2, 5, 8, ... a= -1 given -1+3 =2, 2+3 =5, 5+3=8 and so on so common difference d =3 hence an = -1 +(n-1)*3 @em2000

OpenStudy (anonymous):

thank u so much @princeharryyy

OpenStudy (princeharryyy):

:)